OpenStudy (anonymous):

A Porsche reaches a velocity of 43.0 km/h from a standing start in 15.5 s. What distance does it cover in doing so?

OpenStudy (anonymous):

43 km/hr = 43,000 m / hr = 43,000 / 3600 = 11.944 m /s distance = vel. * time = 11.944 * 15.5 = 185.139 m

OpenStudy (anonymous):

That is incorrect. It is not travelling at constant velocity, hence the equation used by him is wrong. 43 km/h = 11.944m/s v2 = final velocity = 11.944 m/s v1 = initial velocity = 0 m/s (the term "standing start" implies this) t = 15.5s a = (v2 - v1)/t = 0.770581 m/s^2 d = (v1)(t) + 0.5(a)(t)^2 = 0 + 0.5(0.770581)(15.5)^2 = 92.566 m Hence the distance covered is 92.566m.